Email emergency broadcast

                       Tony Finch ⟨fanf2@cam.ac.uk⟩

Mail Support University of Cambridge Computing Service

August 2009 ^{∗}

1 Requirements

• Send email to all @cam email addresses.

• Use of the facility is controlled by a password in the Subject line of the message.

• The message will also be displayed on the Hermes Webmail login page.

• No action is required by the email system managers.

• As simple as possible, to minimise the chance of problems during an emergency.

Note: This is a draft. In particular, the details of liaison between the Computing Service and staff responsible for emergency broadcast are still to be determined. See section 4.

2 Sending an emergency broadcast message

To use the email emergency broadcast system, send a message to ⟨ucam-emergency-broadcast@lists. cam.ac.uk⟩. You must use an email address at cam.ac.uk or at any subdomain of cam.ac.uk.

The Subject line of your message must start with the password in square brackets. The password must match exactly; upper case “A” is considered different from lower case “a”. The following example does not include the real password.

  Subject: [PassWord] evacuate West Cambridge site

Note: The message will be broadcast to all @cam addresses without human intervention. There is no moderation step, and there are no restrictions on who can send a broadcast message, except they must use a valid cam.ac.uk email address and include the correct password in the subject line.

3 Sending a test message

There is a facility for testing all the internal mechanisms of the emergency broadcast system, except for sending the message to all @cam addresses and updating the Hermes Webmail login page. To use it, send a message to ⟨ucam-test-broadcast@lists.cam.ac.uk⟩ with a password in the Subject line as described in the previous section.

∗ $Cambridge:  hermes/doc/misc/emergency-broadcast.tex,v 1.16 2009/08/13 14:37:45 fanf2 Exp $

A test message causes two messages to be sent to the ⟨ucam-broadcasters@lists.cam.ac.uk⟩ mailing list. It modified as described in section 5 then delivered to the list instead of to all @cam addresses. It is further stripped as described in section 6 then sent to the list as a second message instead of being inserted into the webmail login page.

4 Disseminating the password

Note: Details in this section have not been finalized.

The mailing list ⟨ucam-broadcasters@lists.cam.ac.uk⟩ is also intended for liaison between the Computing Service and staff who are responsible for emergency communications, as well as receiving test messages as described in the previous section.

At the moment membership of the list is managed by Computing Service staff. This should be delegated to people more directly responsible for emergency communications.

We suggest that the password is updated on a regular schedule. The mailing list should not be directly used to transmit the password, but it can be used to inform the appropriate people that a new password will be conveyed by other means. The password can only be changed by the email system managers.

5 How messages are altered before broadcast

The body of a message is not changed. Emergency broadcast messages should be short and should not use fancy formatting or attachments. Senders should take care that they use a suitable signature block.

Note: There are no technical mechanisms to enforce these posting guidelines. This is in order to keep the mechanisms as simple as possible.

A number of changes are made to the message headers.

The password is deleted from the Subject line, and “[ EMERGENCY BROADCAST ]” is appended. This tag is added to the end so that as much of the substantive text as possible appears in recipients’ inbox listings. After processing the example looks like:

  Subject: evacuate West Cambridge site  [ EMERGENCY BROADCAST ]

All address lines are deleted (Sender, From, To, CC, BCC, Reply-To) to be replaced by the following lines. The From line is designed so that “EMERGENCY BROADCAST” will appear in recipients’ inbox listings.

  From: EMERGENCY BROADCAST <ucam-emergency-broadcast-do-not-reply@lists.cam.ac.uk>
  To: University of Cambridge Emergency Broadcast :;

6 Updating the Hermes Webmail login page

Every five minutes an automated task checks for emergency broadcast messages and updates the Hermes Webmail login page if necessary. All recent emergency broadcast messages are displayed on the login page, most recent first (the same order as a blog). Messages can only be erased by the email system managers.

Substantial changes are made to the broadcast message. The email headers other than Subject and Date are stripped off and any fancy formatting is removed. Stripping the message avoids any unwanted interactions between webmail’s HTML and any HTML in the broadcast message. Similar changes are not made by the main broadcast mechanism to keep the more important functionality simpler.

7 Handling unwanted messages

Messages sent to ⟨ucam-emergency-broadcast@lists.cam.ac.uk⟩ or ⟨ucam-test-broadcast@lists.

cam.ac.uk⟩ from a non-cam.ac.uk address and/or without the correct password are rejected with an error message saying “Authorization check failed”.

Replies to broadcast messages are rejected with an error message saying “Please do not reply to broadcast messages”. Out-of-office auto-replies and delivery failure notifications are discarded.